OpenStudy (anonymous):

Maximize Q = xy2, where x and y are positive numbers, such that x + y2 = 4.

myininaya (myininaya):

y^2=4-x Q=x(4-x)=4x-x^2 Q'=4-2x set Q'=0 to find critcal numbers 4-2x=0 so x=2 then y^2=4-2=2 so y=sqrt(2) Q=xy^2=2(2)=4